OpenStudy (freckles):

Example for @SolomonZelman \[\int_1^4 \sqrt{x} dx= \frac{2x^\frac{3}{2}}{3} |_1^4 = \\ \frac{2(4)^\frac{3}{2}}{3}-\frac{2(1)^\frac{3}{2}}{3} \\ =\frac{2(2)^3}{3}-\frac{2}{3}=\frac{16}{3}-\frac{2}{3}=\frac{14}{3}\]

OpenStudy (solomonzelman):

Mine is better

OpenStudy (solomonzelman):

\(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~{{\Huge |}_{a}^{b}}}\)

OpenStudy (solomonzelman):

\(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~{{\Huge |}_{a}^{b}}}\) and it is inline

OpenStudy (solomonzelman):

I made it up though,

OpenStudy (freckles):

ok you can show me but I probably won't care to remember :p

OpenStudy (solomonzelman):

`\(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~{{\Huge |}_{a}^{b}}}\)`

OpenStudy (solomonzelman):

I am putting | with a size huge but I am bounding the size Huge like this: { `{`\Huge | `}`_{a}^{b}}

OpenStudy (solomonzelman):

\huge{ } won't bound the size { \huge } will bound the size.

OpenStudy (solomonzelman):

What I was trying to ask thjough...

OpenStudy (solomonzelman):

Is, if there is a natural, designed for this latex, or not?

OpenStudy (freckles):

I have only ever typed F(x)|_a^b

OpenStudy (freckles):

I never tried to look up if there is a "natural way" to type it

OpenStudy (solomonzelman):

it is easy to remember. (Once it is typed from eq. editor INTO the text) 1) you change [ ] to ( ) 2) add " \displaystyle " (to the integral latex, in front) 3) and your " | " thingy, bound with size "\Huge"

OpenStudy (freckles):

http://en.wikibooks.org/wiki/LaTeX/Mathematics

OpenStudy (freckles):

they actually do that same thing here

OpenStudy (freckles):

\left.\frac{x^3}{3}\right|_0^1 \[\left.\frac{x^3}{3}\right|_0^1 \] except they do left and right thingy

OpenStudy (solomonzelman):

I was working on it, but luckily I came up with this, \(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~{{\Huge |}_{a}^{b}}}\)

OpenStudy (solomonzelman):

Wait

OpenStudy (solomonzelman):

yours, \(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~~\left.F(x)\right|_a^b}\) mine \(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~{{\Huge |}_{a}^{b}}}\)

OpenStudy (solomonzelman):

your is high wit a fraction only, that is how "left and right" make it

OpenStudy (solomonzelman):

I can do it though

OpenStudy (freckles):

wikepedia's: \[\int_a^bf(x)dx=\left. F(x) \right|_a^b\] It doesn't look that different from mine

OpenStudy (solomonzelman):

\(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~\left.\color{white}{\frac{.}{.}}\right|_0^1}\)

OpenStudy (solomonzelman):

I am putting a white fraction there in the mid

OpenStudy (freckles):

So I don't understand why the \left. \right it doesn't seem to give it anything extra

OpenStudy (solomonzelman):

\(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~\left.\color{white}{\frac{\frac{\frac{\frac{}{}}{\frac{\frac{}{}}{}}}{\frac{\frac{}{}}{\frac{}{}}}}{\frac{\frac{}{}}{\frac{}{}}}}\right|_a^b}\)

OpenStudy (solomonzelman):

jk

OpenStudy (solomonzelman):

I am putting black white fractions on top of each other.

OpenStudy (solomonzelman):

``` \(\large\color{red}{\displaystyle\int\limits_{a}^{b}f(x)~dx= ~F(x)~{{\Huge |}_{a}^{b}}}\) ``` would be the ideal, for now I guess. (until, and if, I come up with something new)

OpenStudy (solomonzelman):

tnx for putting this question, I guess wiki doesn't know any latex either.

OpenStudy (freckles):

lt looks like it knows some things

OpenStudy (solomonzelman):

does, but not enough I guess. I outsmarted it:)

OpenStudy (solomonzelman):

@ganeshie @iambatman @Kainui Chance 4 u to see. (whenever you have time) Check it out. (latex) \(\large\color{green}{ \displaystyle \int\limits_{a}^{b}f(x)~dx= ~F(x)~{ {\Huge |}_{a}^{b} } }\) :D \(\large\color{blue}{ \Downarrow }\) ``` \(\large\color{green}{ \displaystyle \int\limits_{a}^{b}f(x)~dx= ~F(x)~{ {\Huge |}_{a}^{b} } }\) ``` \(\large\color{blue}{ \Uparrow }\) (bounded the | with { } )

OpenStudy (solomonzelman):

@ganeshie8 (failed to tag u, sorry)

ganeshie8 (ganeshie8):

Ahh thats useful, i tried to stop `\Huge` earlier but failed, i know how to stop it now thanks :)

ganeshie8 (ganeshie8):

i use below usually ` \left. \dfrac{x^2}{2} \right|_a^b` produces below \[ \left.\dfrac{x^2}{2}\right|_a^b\] ` \right| ` sees stuff between the bars and scales the vertical bar accordingly